how long will it take for 500 to double if it is invested at 6% anniual interest compouded continuously
Answer by ewatrrr(24785)

A = P e ^(rt),
A/P = e ^(rt)
.......
2 = e ^(.06t)
ln(2)/.06 = t = 11.5525 yrs